A pharmacist is to prepare 20 ml of a special eye drop for a glaucoma patient the eye drop solution must have a 3 % active ingredient but the pharmacist only has 15% and 2% in stock how much of each type of solution should be used to fill the prescription?
Answer by Alan3354(69443)

A pharmacist is to prepare 20 ml of a special eye drop for a glaucoma patient the eye drop solution must have a 3 % active ingredient but the pharmacist only has 15% and 2% in stock how much of each type of solution should be used to fill the prescription?
---------------
f = amount of 15%
t = amount of 2%
-----
f + t = 20 (total solution)
15f + 2t = 3*20 (total active ingredient)
----------
solve for f & t